what is a step by step process to find radius for 81pi ft squared?

Answer by stanbon(75887)

what is a step by step process to find radius for 81pi ft squared?
----
Assuming you are referring to the radius of a circle.
Area = pi*r^2
------
Equation:
pi*r^2 = 81*pi
-------
Divide by pi to get:
r^2 = 81
r = 9 ft.
===============
Cheers,
Stan H.